Circular No. 7967

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002ez, 2002fa, AND 2002fb Three new supernovae have been discovered on unfiltered CCD images,

as listed below. SNe 2002ez and 2002fa were reported by W. M. Wood-Vasey et al. (cf. IAUC 7964, to which should be added B. Farris) on NEAT images obtained at Palomar, while SN 2002fb was found by LOTOSS (cf. IAUC 7906) on KAIT images and reported by D. Weisz and W. D. Li.

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002ez July 23.28 22^{h}00^{m}20^{s}.93 −14^{o}00′22′′.5 17.6 4′′ E, 18′′ S
2002fa Aug. 24.34 20 52 21.78 + 2 08 41.9 19.4 4′′.1 E
2002fb Sept. 6.4 1 57 48.90 +36 20 26.3 17.1: 17′′.5 W, 9′′.1 S

Additional magnitude estimates from the same respective observatories for each object: SN 2002ez, 2001 Sept. 6 UT, [20.5; Nov. 27, [20.5; 2002 July 23.40, 17.6; 29, 18.5; Aug. 8, 18.4; 20, 18.9; 26, 19.3. SN 2002fa, July 1, [20.5; 20, [20.5; Aug. 3, [20.5; Aug. 24.39, 19.4; Sept. 1, 19.3. SN 2002fb in NGC 759, Aug. 31.4, [19.0:; Sept. 7.4, 16.9: (1-m Nickel telescope).

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002fb, obtained by M. Calkins on Sept. 7.50 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a possibly peculiar type-Ia supernova near maximum. Several titanium features appear stronger than normal, implying that this may be a subluminous type-Ia event similar to SN 1991bg or SN 1999by. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4667 km/s for the host galaxy, the supernova expansion velocity is ∼ 11600 km/s for Si II (rest 635.5 nm). The spectralfeature age of the supernova is 2 ± 2 days before maximum light.

SUPERNOVA 2002et IN MCG –04-47-10 A. Clocchiatti and P. Lagos, Pontificia Universidad Católica de Chile;

and S. Benetti, A. Pastorello, and M. Riello, Osservatorio Astronomico di Padova, report that a spectrum (range 346–958 nm, resolution 0.5 nm) obtained on Aug. 31.10 UT with the European Southern Observatory 3.6-m telescope (+ EFOSC2) indicates that SN 2002et (cf. IAUC 7959) is a type-Ia supernova. The best match is obtained with a spectrum of SN 1989B, ≈ 10 days after maximum light. The redshift of the parent galaxy (measured from weak stellar Na D lines in the parent galaxy at 605.4 nm and Hα in emission at 673.7 nm) is 0.0273.

2002 September 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green